OpenStudy (anonymous):

Approximate the real number solution(s) to the polynomial function f(x) = x3 + 2x2 - 5x - 6

OpenStudy (anonymous):

answers: x = -3, x = -1, x = 2 x = 3, x = 1, x = -2 x = 3, x = -1, x = 2 x = -3, x = 1, x = -2

OpenStudy (anonymous):

=x^3+x^2+x^2+x-6x-6 =x^2(x+1)+x(x+1)-6(x+1) =(x+1)(x^2+x-6) thus x=-1,x=-3,x=2